Three equal charges q are fixed at the corners of
Answer #1

2B)

v = k(q/a + q/a + q/a)

v = 3kq / a

where, k = 1/4pi*eo

v = 3q / 4pi*eo*a

W = qv

W = 3q^2 / 4*pi*eo*a

Ans(X) - 3q^2 / 4*pi*eo*a

2C)

Ans = q^2 / 4*pi*eo*a